OpenStudy (anonymous):

Us Quotient rule to differentiate the function f(x)=3x/(x^(4)+8)

OpenStudy (ybarrap):

d/dx((3 x)/(x^4+8)) Factor out constants: = 3 (d/dx(x/(x^4+8))) Use the quotient rule, d/dx(u/v) = (v ( du)/( dx)-u ( dv)/( dx))/v^2, where u = x and v = x^4+8: = 3 ((x^4+8) (d/dx(x))-x (d/dx(x^4+8)))/(x^4+8)^2 The derivative of x is 1: = (3 (1 (x^4+8)-x (d/dx(x^4+8))))/(x^4+8)^2 Differentiate the sum term by term: = (3 (-x (d/dx(x^4)+d/dx(8))+x^4+8))/(x^4+8)^2 The derivative of 8 is zero: = (3 (-x (d/dx(x^4)+0)+x^4+8))/(x^4+8)^2 The derivative of x^4 is 4 x^3: = (3 (x^4-(4 x^3) x+8))/(x^4+8)^2